# Axionlike particle-assisted supercooling chiral phase transition in QCD: Identifying Coleman-Weinberg type-chiral phase transition in QCD-like scenarios

### v2

@@ II Two-flavor mean-field NJL monitor
where we have taken the isospin symmetric limit: $m_{u}=m_{d}=m$, and have also introduced the determinant term related to the $U(1)_{A}$ anomaly in the underlying QCD; $\tau_{a}$ denotes the Pauli matrices with $\tau_{0}=1_{2\times 2}$. The current quark mass is defined to be real and positive, so that the strong CP phase $\theta$ coupled to the $U(1)$ axial anomalous determinant term has already involved the quark mass phase, i.e., $\theta\equiv\bar{\theta}$. Regarding the scale violation classes at this point, as listed in the Introduction, we see that
@@
We work in the MFA and take into account the mean field only for the $SU(2)$ isospin singlet #1#1#1This assumption is ensured by absence of the isospin chemical potential. Our current discussions can also be straightforwardly extended by including the isospin breaking.. Then the targeted mean fields are those for the following bilinears: $\left(\bar{q}^{\prime}_{i}q^{\prime}_{i}\right)$ and $\left(\bar{q}^{\prime}_{i}i\gamma_{5}q^{\prime}_{i}\right)$. Those are expanded around the mean fields $S^{\prime}=\langle\bar{q}^{\prime}_{i}q^{\prime}_{i}\rangle$ and $P^{\prime}=\langle\bar{q}^{\prime}_{i}i\gamma_{5}q^{\prime}_{i}\rangle$ as $\bar{q}^{\prime}_{i}q^{\prime}_{i}=S^{\prime}+\left(:\bar{q}^{\prime}_{i}q^{\prime}_{i}:\right)$ and $\langle\bar{q}^{\prime}_{i}i\gamma_{5}q^{\prime}_{i}\rangle=P^{\prime}+\left(:\bar{q}^{\prime}_{i}i\gamma_{5}q^{\prime}_{i}:\right)$ #2#2#2The U(1) axial rotation relates the scalar and pseudoscalar bilinears between the original- and prime-base scalar and pseudoscalar bilinears (for each quark flavor $i$) as $\left(\bar{q}_{i}q_{i}\right)$ $=\left(\bar{q}^{\prime}_{i}q^{\prime}_{i}\right)\text{cos}\frac{\theta}{2}+\left(\bar{q}^{\prime}_{i}i\gamma_{5}q^{\prime}_{i}\right)\text{sin}\frac{\theta}{2}\,$, $\left(\bar{q}_{i}i\gamma_{5}q_{i}\right)$ $=-\left(\bar{q}^{\prime}_{i}q^{\prime}_{i}\right)\text{sin}\frac{\theta}{2}+\left(\bar{q}^{\prime}_{i}i\gamma_{5}q^{\prime}_{i}\right)\text{cos}\frac{\theta}{2}\,.$ . Here the terms sandwiched by ”:” stand for the normal order product, which means $\langle:\mathcal{O}:\rangle=0$ for $\mathcal{O}=S^{\prime},P^{\prime}$, hence ${\cal L}={\cal L}|_{\rm MF-only}+(:{\cal L}:)$. Thus, in the MFA, the Lagrangian in Eq.(II) takes the form
@@
As is evident in the form of Eq.(6), $\mu_{5}$ and $\theta$ merely act as a catalyzer against the CW-type scale anomaly, i.e., the driving force for the phase transition in the case without them. For instance, even for a large $\mu_{5}$ at $T=\mu_{B}=0$, in the case with massive quarks including physical point, $\mu_{5}$ does not generate the Debye mass or logarithmic term for $M$, hence $\mu_{5}$ itself does not trigger the chiral restoration #3#3#3In the chiral limit, $\mu_{5}$ can act like $\mu_{B}$ as seen from Eq.(7), and therefore can trigger the first order phase transition as has been discussed in the literature Chernodub:2011fr . Even at physical point, a dynamically induced $\mu_{5}$ as the mean field of the axialvector field can also serve as the source for triggering the first order phase transition Shao:2022oqw , which we will not consider in the present study. In the present work, we will not consider this possibility just for simplicity.. The MFA-NJL with nonzero $\mu_{5}$, however, involves the high regularization scheme dependence Yu:2015hym . In the literature Farias:2016let a proper subtraction scheme has been shown to support the chiral crossover, not the first order, even in the chiral limit, where $\mu_{5}$ acts as a catalyzer to enhance the chiral/axial breaking via efficiently driving the topological charge fluctuation to the axial anomaly. We will come back to this point later on.
@@ III Exploring CW-type chiral phase transition
In this section, we discuss the features of the chiral phase transition based on the thermodynamic potential in Eq.(6) in light of realization of the CW-type phase transition on the basis of the scale violation classification. The cases we shall study are classified into two: with $\mu_{B}=0$ (Sec. III.1) or $\mu_{B}\neq 0$ (Sec. III.2). Our main phenomenological proposal is in the former case, where an ALP is predicted to deform the conventional chiral crossover/phase transition into the CW-type supercooling, which we call the ALP-assisted CW-type chiral phase transition. This particular scenario will separately be addressed in more detail, in the later section, Sec. IV.
@@ III.1 The case with $\mu_{B}=0$
It is well known in the two-flavor NJL model that at the physical point for quark masses, the chiral phase transition at vanishing baryon chemical potentials is crossover (see, e.g., a review Hatsuda:1994pi ). This is also true even with nonzero $\theta$, as long as $\theta\neq\pi$, at which case the CP phase transition along $P=S^{\prime}$ undergoes like second order Boer:2008ct ; Boomsma:2009eh ; Sakai:2011gs ; Sasaki:2013ewa ; Lu:2018ukl ; Huang:2024nbd . Now we understand this phase transition property in terms of the scale violation classes.
First of all, we take $\theta=0$ just for simplicity. It turns out that this will not essentially lose generality on the phase transition property even with $\theta\neq 0$, as long as $0<\theta<\pi$. With $\theta=0$, we have $S^{\prime}=S$ and $P^{\prime}=P=0$, or equivalently, ($\beta=0$ in Eq.(II), hence $M^{2}=\alpha^{2}$ in Eq.(7). In that case Eq.(8) can still further be simplified to
@@
Consider first the vacuum case with $T=0$. The tree-level (or bare) mass term, $(M-m)^{2}$, has come from the four-fermion interactions along with $g_{s}$ and $g_{d}$, including the shift by the current quark mass. This is a soft-scale breaking term. In addition, the first term in the loop correction part (inside the square bracket in Eq.(9)) includes quadratic $\sim M^{2}\Lambda^{2}$ and logarithmic divergent terms $\sim M^{4}\ln(\Lambda/M)$, where $\Lambda$ denotes the loop momentum cutoff. The former definitely takes negative sign, and gives a destructive contribution to the bare $M^{2}$ term, which destabilizes the chiral symmetric phase at $M=0$ up to the linear shift by $m$, hence triggers the dynamical chiral symmetry breaking. This is another soft-scale breaking term. The net soft-breaking term takes largely negative sign at $T=0$. Meanwhile, the latter logarithmic divergent term takes nothing but a form of the CW-type scale anomaly. Thus, at $T=0$, the effective potential derived from Eq.(9) (with renormalization prescribed) can be cast into essentially the same form as in the generic CW-type potential given in Eq.(1), with a large negative curvature at around $M=0$ ($m_{0}^{2}<0$ in Eq.(1)):
@@
At finite $T$, the negative curvature at round $M=0$ gets thermal corrections (the second term in square bracket in Eq.(9)). This starts to act as a cancellation of the soft-scale breaking term $\propto M^{2}$. At some high enough $T$, the original soft-breaking term, i.e. the dynamical chiral symmetry breaking will be canceled (i.e., $m_{0}^{2}=0$ in Eq.(1)), that determines the typical critical temperature $T_{c}$. As $T$ gets further higher, the soft-breaking term is instead developed by the thermal mass term $\sim T^{2}M^{2}$. At this moment, one might suspect that a potential barrier is built around $M=0$ by the positive curvature (the thermal mass) $\sim(+T^{2}M^{2})$ and the logarithmic damping term $\sim(-M^{4}\ln M)$, just like the case of the CW-type potential at high $T$, hence the thermal chiral phase transition goes like first order or more like a supercooling form. However, it is not the case because of the nontrivial cancellation for the original CW-type logarithmic term by the high $T$ correction.
To see this cancellation more explicitly, we work on the high $T$ expansion for the thermal correction term in Eq.(9), which goes like
$\left.-N_{c}N_{f}\cdot T\int\frac{d^{3}p}{(2\pi)^{3}}\ln(1+e^{-\frac{E_{0}}{T}})\right|_{T\gg M}$ $=N_{c}N_{f}\left[-\frac{7\pi^{2}}{180}T^{4}+\frac{1}{12}T^{2}M^{2}+\frac{1}{16\pi^{2}}M^{4}\ln M^{2}-\frac{1}{16\pi^{2}}M^{4}\ln a_{f}T^{2}+\cdots\right]\,$, (11)
where $a_{f}=2\ln\pi-2\gamma_{E}\simeq 1.14$. As seen from Eqs.(10) and (11), the original CW-type logarithmic term $\sim M^{4}\ln M$ is precisely canceled each other. Thereby, the CW-type first-order phase transition is not realized: no barrier is generated. This is completely consistent with the well-known fact that the two-flavor chiral phase transition in the MFA-NJL model, with $\mu_{B}=0$, is second order in the chiral limit with $m=0$, and crossover at physical point. This phenomenon can also be understood in a view of the original scale anomaly cancellation in an extremely hot medium, which implies a free quark gas picture simply following the Stephan-Boltzmann form of pressure, $\Omega\sim p\sim T^{4}$ (corresponding to the first term in Eq.(11)). This cancellation is further promoted to work by nonzero $\mu_{5}$, because $\mu_{5}$ can act as a catalyzer for the chiral/axial breaking when a proper regularization scheme is carried out Farias:2016let .
@@
This scale anomaly cancellation does not happen when the soft-scale breaking mass is small enough as in the case that has been addressed in the literature Gupta:2011ez . This is because the high $T$ expansion will not well approximate the thermal potential due to the scale invariance at around $M=0$, hence a barrier is created and the phase transition undergoes like first order and supercooling. Thus, the scale anomaly cancellation is intrinsic to a negatively large soft-scale breaking case, which makes the curvature of the potential deeply negative at around $M=0$, like in the present NJL scenario with the significant dynamical chiral symmetry breaking.
@@
Even when nonzero $\theta$ is turned on, we can again arrange $\Omega$ so that the thermal loop and vacuum-one loop terms are still singly dependent of $M$. Nonzero $\theta$ effects only come into the tree-level part, in which either $P^{\prime}=\frac{m\sin\frac{\theta}{2}-\beta}{2(g_{s}-g_{d})}$ or $S^{\prime}=\frac{m\cos\frac{\theta}{2}-\alpha}{2(g_{s}+g_{d})}$ direction can be chosen as the actual mean field directions in addition to the $M=\sqrt{\alpha^{2}+\beta^{2}}$ direction. Thus, the dynamical chiral symmetry breaking is still monitored along the $M$ direction, with keeping the size of the soft-scale symmetry breaking at each $T$. Hence the high $T$ expansion works as in the case with $\theta=0$ and the scale anomaly cancellation is still operative, so the chiral phase transition substantially follows the same way and no first order phase transition can be seen even with $\theta\neq 0$. On the other hand, the $CP$ phase transition feature observed along $S^{\prime}$ or $\alpha$ highly depends on $\theta$, which exhibits a second order type when $\theta=\pi$ Boer:2008ct ; Boomsma:2009eh ; Sakai:2011gs ; Huang:2024nbd .
@@
The dynamical chiral symmetry breaking occurs when the gap equation, i.e., the stationary condition for Eq.(10) gets nontrivial solution. It is the case when the following condition is met:
@@
in the chiral limit, for simplicity. At the criticality where $\frac{N_{f}N_{c}(g_{s}+g_{d})\Lambda^{2}}{2\pi^{2}}=1$, the soft-scale breaking mass becomes zero: $m_{0}^{2}=0$ in terms of Eq.(1). In two-flavor QCD case, we have $\frac{N_{f}N_{c}(g_{s}+g_{d})\Lambda^{2}}{2\pi^{2}}\sim 1.5$, which is fixed by fitting to QCD hadron observables (see also Eq.(15)). Thus, QCD indeed yields a significantly large soft-scale breaking as dynamical chiral symmetry breaking. Note that the near criticality condition, $\frac{N_{f}N_{c}(g_{s}+g_{d})\Lambda^{2}}{2\pi^{2}}\sim 1$, also implies the dynamics near an ultraviolet fixed point (in the chiral broken phase for $\frac{N_{f}N_{c}(g_{s}+g_{d})\Lambda^{2}}{2\pi^{2}}>1$), $\Lambda\frac{\partial[(g_{s}+g_{d})\Lambda^{2}]}{\partial\Lambda}\sim 0$, i.e., the almost quantum scale invariance.
In the MFA, the near criticality condition for the soft-scale breaking, $\frac{N_{f}N_{c}(g_{s}+g_{d})\Lambda^{2}}{2\pi^{2}}=1$, implies
@@
where we have used Eq.(II) with $\theta=m=0$. Two-flavor QCD gives $\frac{M\Lambda^{2}}{\langle-\bar{q}q\rangle/N_{f}}\sim 20$. For a fixed place of $S=\langle\bar{q}q\rangle$ in the thermodynamic potential, Eq.(13) requires a smaller dynamical (full) mass $M$ by about 40%, compared to the existing two-flavor QCD. Since generically $\langle-\bar{q}q\rangle\sim M^{3}$ (when evaluated at the renormalization scale $\mu=M$), one may read the condition in Eq.(13) as $\Lambda/M\gg 1$. This indicates a large scale hierarchy between the ultraviolet and infrared scales intrinsic to the underlying theory contrast to QCD: ordinary QCD yields almost the same order for them, $\Lambda\sim(2-3)M$. The soft-scale breaking is ensured by the nature of the almost scale-invariant gauge theory, and the CW-type scale anomaly is induced by nonperturbative running of the gauge coupling associated with the dynamical chiral symmetry breaking Leung:1985sn . One candidate theory to realize such a large scale gap is many flavor QCD, e.g., QCD with $N_{f}=8$ Aoki:2013xza ; LSD:2014nmn ; Hasenfratz:2014rna . In this scenario, the approximate quantum scale symmetry gets anomalous dominantly by the CW-type scale anomaly of the form $\sim M^{4}\ln M$. This characteristic feature has so far been applied to various phenomena, such as dynamical CW inflation scenarios Ishida:2019wkd ; Cacciapaglia:2023kat ; Zhang:2023acu ; Zhang:2024vpp ; Liu:2024xrh ; Zhang:2025kbu ; Cacciapaglia:2025xqd , models of dynamical electroweak symmetry breaking Matsuzaki:2012gd ; Matsuzaki:2012vc ; Matsuzaki:2012mk ; Matsuzaki:2013eva ; Matsuzaki:2015sya ; Kasai:2016ifi ; Hansen:2016fri ; Appelquist:2017wcg ; Appelquist:2017vyy ; Cata:2019edh ; Appelquist:2019lgk ; Brown:2019ipr ; Elander:2025fpk ; Faedo:2024zib ; Appelquist:2024koa ; Ingoldby:2023mtf ; Freeman:2023ket ; LSD:2023uzj ; Choi:2011fy ; Choi:2012kx ; Appelquist:2022mjb ; Appelquist:2022qgl ; Golterman:2020utm ; Miura:2018dsy , and phenomenological applications to QCD dilaton Kawaguchi:2020kce ; Crewther:2013vea ; Li:2016uzn ; Fujii:2025aip ; Sheng:2025jvn ; Zhang:2024iye ; Zhang:2024sju ; Bersini:2023uat ; Ma:2023ugl ; Shao:2022njr ; Cata:2019edh ; Li:2017udr ; Paeng:2017qvp ; Ma:2016gdd ; Ma:2016nki .
Below we will propose a new scenario at $\mu_{B}\sim 0$, which realizes the approximate scale invariance only at around $T=T_{c}$ (Sec. IV). Before this proposal, in the next section we shall take a look at the case with nonzero $\mu_{B}$, where the realization of the supercooling chiral phase transition, as has been discussed in the literature Shao:2024dxt ; Shao:2024ygm , will be recapped in terms of the class of the soft-scale breaking and the Debye mass effect.

### v4

@@ II Two-flavor mean-field NJL monitor
where we have taken the isospin symmetric limit: $m_{u}=m_{d}=m$, and have also introduced the determinant term related to the $U(1)_{A}$ anomaly in the underlying QCD; $\tau_{a}$ denotes the Pauli matrices with $\tau_{0}=1_{2\times 2}$. The current quark mass is defined to be real and positive, so that the strong CP phase $\theta$ coupled to the $U(1)$ axial-anomalous determinant term has already involved the quark mass phase, i.e., $\theta\equiv\bar{\theta}$. Regarding the scale violation classes at this point, as listed in the Introduction, we see that
@@
We work in the MFA and take into account the mean fields only for the $SU(2)$ isospin singlets #1#1#1This assumption is ensured by the absence of the isospin chemical potential, and our current discussions can also be straightforwardly extended by including the isospin breaking.. Then the target mean fields are those for the following bilinears: $\left(\bar{q}^{\prime}_{i}q^{\prime}_{i}\right)$ and $\left(\bar{q}^{\prime}_{i}i\gamma_{5}q^{\prime}_{i}\right)$. Those are expanded around the mean fields $S^{\prime}=\langle\bar{q}^{\prime}_{i}q^{\prime}_{i}\rangle$ and $P^{\prime}=\langle\bar{q}^{\prime}_{i}i\gamma_{5}q^{\prime}_{i}\rangle$ as $\bar{q}^{\prime}_{i}q^{\prime}_{i}=S^{\prime}+\left(:\bar{q}^{\prime}_{i}q^{\prime}_{i}:\right)$ and $\langle\bar{q}^{\prime}_{i}i\gamma_{5}q^{\prime}_{i}\rangle=P^{\prime}+\left(:\bar{q}^{\prime}_{i}i\gamma_{5}q^{\prime}_{i}:\right)$ #2#2#2The U(1) axial rotation relates the scalar and pseudoscalar bilinears between the original- and prime-base scalar and pseudoscalar bilinears (for each quark flavor $i$) as $\left(\bar{q}_{i}q_{i}\right)$ $=\left(\bar{q}^{\prime}_{i}q^{\prime}_{i}\right)\text{cos}\frac{\theta}{2}+\left(\bar{q}^{\prime}_{i}i\gamma_{5}q^{\prime}_{i}\right)\text{sin}\frac{\theta}{2}\,$, $\left(\bar{q}_{i}i\gamma_{5}q_{i}\right)$ $=-\left(\bar{q}^{\prime}_{i}q^{\prime}_{i}\right)\text{sin}\frac{\theta}{2}+\left(\bar{q}^{\prime}_{i}i\gamma_{5}q^{\prime}_{i}\right)\text{cos}\frac{\theta}{2}\,.$ . Here the terms sandwiched by ”:” stand for the normal order product, which means $\langle:\mathcal{O}:\rangle=0$ for $\mathcal{O}=S^{\prime},P^{\prime}$, hence ${\cal L}={\cal L}|_{\rm MF-only}+(:{\cal L}:)$. Thus, in the MFA, the Lagrangian in Eq.(II) takes the form
@@
As is evident in the form of Eq.(6), $\mu_{5}$ and $\theta$ merely act as a catalyzer against the CW-type scale anomaly, i.e., the driving force for the phase transition in the case without them. For instance, even for a large $\mu_{5}$ at $T=\mu_{B}=0$, in the case with massive quarks including physical point, $\mu_{5}$ does not generate the Debye mass or logarithmic term for $M$, hence $\mu_{5}$ itself does not trigger the chiral restoration #3#3#3In the chiral limit, $\mu_{5}$ can act like $\mu_{B}$ as seen from Eq.(7), and therefore can trigger the first order phase transition as has been discussed in the literature Chernodub:2011fr . Even at physical point, a dynamically induced $\mu_{5}$ as the mean field of the axialvector field can also serve as the source for triggering the first order phase transition Shao:2022oqw , which we will not consider in the present study. . The NJL-MFA with nonzero $\mu_{5}$, however, involves the high regularization scheme dependence Yu:2015hym . In the literature Farias:2016let a proper subtraction scheme has been shown to support the chiral crossover, not the first order, even in the chiral limit, where $\mu_{5}$ acts as a catalyzer to enhance the chiral/axial breaking via efficiently driving the topological charge fluctuation to the axial anomaly. We will come back to this point later on.
@@ III Exploring CW-type chiral phase transition
In this section, we discuss the features of the chiral phase transition based on the thermodynamic potential in Eq.(6) in light of realization of the CW-type phase transition on the basis of the scale violation classification. The cases we shall study are classified into two: with $\mu_{B}=0$ (Sec. III.1) or $\mu_{B}\neq 0$ (Sec. III.2). Our main phenomenological proposal is in the former case, where an ALP is predicted to deform the conventional chiral crossover/phase transition into the CW-type supercooling, which we call the ALP-assisted CW-type chiral phase transition. This particular scenario will separately be addressed in more detail in the later section, Sec. IV.
@@ III.1 The case with $\mu_{B}=0$
It is well known in the two-flavor NJL model that at the physical point for quark masses, the chiral phase transition at vanishing baryon chemical potentials is crossover (see, e.g., a review Hatsuda:1994pi ). This is also true even with nonzero $\theta$, as long as $\theta\neq\pi$, in which case the CP phase transition along $P=S^{\prime}$ undergoes like a second order Boer:2008ct ; Boomsma:2009eh ; Sakai:2011gs ; Sasaki:2013ewa ; Lu:2018ukl ; Huang:2024nbd . Now we understand this phase transition property in terms of the scale violation classes.
First, we take $\theta=0$ for simplicity. It turns out that this will not essentially lose generality on the phase transition property even with $\theta\neq 0$, as long as $0<\theta<\pi$. With $\theta=0$, we have $S^{\prime}=S$ and $P^{\prime}=P=0$, or equivalently, $\beta=0$ in Eq.(II), hence $M^{2}=\alpha^{2}$ in Eq.(7). In that case Eq.(8) can be simplified further to
@@
Consider first the vacuum case with $T=0$. The tree-level (or bare) mass term, $(M-m)^{2}$, has come from the four-fermion interactions along with $g_{s}$ and $g_{d}$, including the shift by the current quark mass. This is a soft-scale breaking term. In addition, the first term in the loop correction part (inside the square brackets in Eq.(9)) includes quadratic $\sim M^{2}\Lambda^{2}$ and logarithmic divergent terms $\sim M^{4}\ln(\Lambda/M)$, where $\Lambda$ denotes the loop momentum cutoff. The former definitely takes a negative sign and gives a destructive contribution to the bare $M^{2}$ term, which destabilizes the chiral symmetric phase at $M=0$ up to the linear shift by $m$, hence triggers the dynamical chiral symmetry breaking. This is another soft-scale breaking term. The net soft-breaking term takes a largely negative sign at $T=0$. Meanwhile, the latter logarithmic divergent term takes nothing but a form of the CW-type scale anomaly. Thus, at $T=0$, the effective potential derived from Eq.(9) (with renormalization prescribed) can be cast into essentially the same form as the generic CW-type potential given in Eq.(1), with a large negative curvature around $M=0$ ($m_{0}^{2}<0$ in Eq.(1)):
@@
At finite $T$, the negative curvature around $M=0$ gets thermal corrections (the second term in the square brackets in Eq.(9)). This starts to act as a cancellation of the soft-scale breaking term $\propto M^{2}$. At some high enough $T$, the original soft-breaking term, i.e. the dynamical chiral symmetry breaking will be canceled (i.e., $m_{0}^{2}=0$ in Eq.(1)), which determines the typical critical temperature $T_{c}$. As $T$ gets further higher, the soft-breaking term is instead developed by the thermal mass term $\sim T^{2}M^{2}$. At this moment, one might suspect that a potential barrier is built around $M=0$ by the positive curvature (the thermal mass) $\sim(+T^{2}M^{2})$. In addition, the original CW-type logarithmic term is actually canceled by the high $T$ correction. To see this cancellation more explicitly, we work on the high $T$ expansion for the thermal correction term in Eq.(9), which goes like
$\left.-2N_{c}N_{f}\cdot 2T\int\frac{d^{3}p}{(2\pi)^{3}}\ln(1+e^{-\frac{E_{0}}{T}})\right|_{T\gg M}$ $=N_{c}N_{f}\left[-\frac{7\pi^{2}}{180}T^{4}+\frac{1}{12}T^{2}M^{2}+\frac{1}{16\pi^{2}}M^{4}\ln M^{2}-\frac{1}{16\pi^{2}}M^{4}\ln a_{f}T^{2}+\cdots\right]\,$, (11)
where $a_{f}=2\ln\pi-2\gamma_{E}\simeq 1.14$. As seen from Eq.(10), the original CW-type logarithmic term $\sim M^{4}\ln M$ is precisely canceled.
@@
This scale anomaly cancellation does not happen when the soft-scale breaking mass is small enough as in the case that has been addressed in the literature Gupta:2011ez . This is because the high $T$ expansion will not accurately approximate the thermal potential due to the scale invariance around $M=0$. Thus, the scale anomaly cancellation is intrinsic to a negatively large soft-scale breaking case, which makes the curvature of the potential deeply negative around $M=0$, like in the present NJL scenario with significant dynamical chiral symmetry breaking.
@@
Even when nonzero $\theta$ is turned on, we can again arrange $\Omega$ so that the thermal loop and vacuum-one loop terms are still singly dependent of $M$. Nonzero $\theta$ effects only come into the tree-level part, in which either $P^{\prime}=\frac{m\sin\frac{\theta}{2}-\beta}{2(g_{s}-g_{d})}$ or $S^{\prime}=\frac{m\cos\frac{\theta}{2}-\alpha}{2(g_{s}+g_{d})}$ direction can be chosen as the actual mean field directions in addition to the $M=\sqrt{\alpha^{2}+\beta^{2}}$ direction. Thus, dynamical chiral symmetry breaking is still monitored along the $M$ direction, keeping the size of the soft-scale symmetry breaking at each $T$. Hence the high $T$ expansion works as in the case with $\theta=0$ and the scale anomaly cancellation is still operative, so the chiral phase transition substantially follows the same way. On the other hand, the $CP$ phase transition feature observed along $S^{\prime}$ or $\alpha$ highly depends on $\theta$, which exhibits a second order type when $\theta=\pi$ Boer:2008ct ; Boomsma:2009eh ; Sakai:2011gs ; Huang:2024nbd .
@@
The dynamical chiral symmetry breaking occurs when the gap equation, i.e., the stationary condition for Eq.(10) gets nontrivial solution. It is the case where the following condition is met:
@@
in the chiral limit, for simplicity. At the criticality where $\frac{N_{f}N_{c}(g_{s}+g_{d})\Lambda^{2}}{2\pi^{2}}=1$, the soft-scale breaking mass becomes zero: $m_{0}^{2}=0$ in terms of Eq.(1). In the two-flavor QCD case, we have $\frac{N_{f}N_{c}(g_{s}+g_{d})\Lambda^{2}}{2\pi^{2}}\sim 1.5$, which is fixed by fitting to QCD hadron observables (see also Eq.(15)). Thus, QCD indeed yields a significantly large soft-scale breaking as dynamical chiral symmetry breaking. Note that the near criticality condition, $\frac{N_{f}N_{c}(g_{s}+g_{d})\Lambda^{2}}{2\pi^{2}}\sim 1$, also implies the dynamics near an ultraviolet fixed point (in the chiral broken phase for $\frac{N_{f}N_{c}(g_{s}+g_{d})\Lambda^{2}}{2\pi^{2}}>1$), $\Lambda\frac{\partial[(g_{s}+g_{d})\Lambda^{2}]}{\partial\Lambda}\sim 0$, i.e., the almost quantum scale invariance.
In the MFA, the near criticality condition for soft-scale breaking, $\frac{N_{f}N_{c}(g_{s}+g_{d})\Lambda^{2}}{2\pi^{2}}=1$, implies
@@
where we have used Eq.(II) with $\theta=m=0$. Two-flavor QCD gives $\frac{M\Lambda^{2}}{\langle-\bar{q}q\rangle/N_{f}}\sim 20$. For a fixed place of $S=\langle\bar{q}q\rangle$ in the thermodynamic potential, Eq.(13) requires a smaller dynamical (full) mass $M$ by about 40%, compared to the existing two-flavor QCD case. Since generically $\langle-\bar{q}q\rangle\sim M^{3}$ (when evaluated at the renormalization scale $\mu=M$), the condition in Eq.(13) may be read as $\Lambda/M\gg 1$. This indicates a large scale hierarchy between the ultraviolet and infrared scales intrinsic to the underlying theory contrast to QCD: ordinary QCD yields almost the same order for them, $\Lambda\sim(2-3)M$. Soft-scale breaking is ensured by the nature of the almost scale-invariant gauge theory, and the CW-type scale anomaly is induced by nonperturbative running of the gauge coupling associated with dynamical chiral symmetry breaking Leung:1985sn . One candidate theory to realize such a large scale gap is many flavor QCD, e.g., QCD with $N_{f}=8$ Aoki:2013xza ; LSD:2014nmn ; Hasenfratz:2014rna . In this scenario, the approximate quantum scale symmetry gets anomalous dominantly by the CW-type scale anomaly of the form $\sim M^{4}\ln M$. This characteristic feature has so far been applied to various phenomena, such as dynamical CW inflation scenarios Ishida:2019wkd ; Cacciapaglia:2023kat ; Zhang:2023acu ; Zhang:2024vpp ; Liu:2024xrh ; Zhang:2025kbu ; Cacciapaglia:2025xqd , models of dynamical electroweak symmetry breaking Matsuzaki:2012gd ; Matsuzaki:2012vc ; Matsuzaki:2012mk ; Matsuzaki:2013eva ; Matsuzaki:2015sya ; Kasai:2016ifi ; Hansen:2016fri ; Appelquist:2017wcg ; Appelquist:2017vyy ; Cata:2019edh ; Appelquist:2019lgk ; Brown:2019ipr ; Elander:2025fpk ; Faedo:2024zib ; Appelquist:2024koa ; Ingoldby:2023mtf ; Freeman:2023ket ; LSD:2023uzj ; Choi:2011fy ; Choi:2012kx ; Appelquist:2022mjb ; Appelquist:2022qgl ; Golterman:2020utm ; Miura:2018dsy , and phenomenological applications to QCD dilaton Kawaguchi:2020kce ; Crewther:2013vea ; Li:2016uzn ; Fujii:2025aip ; Sheng:2025jvn ; Zhang:2024iye ; Zhang:2024sju ; Bersini:2023uat ; Ma:2023ugl ; Shao:2022njr ; Cata:2019edh ; Li:2017udr ; Paeng:2017qvp ; Ma:2016gdd ; Ma:2016nki .
In the following, we will propose a new scenario for $\mu_{B}\sim 0$, which realizes the approximate scale invariance only at around $T=T_{c}$ (Sec. IV). Before this proposal, in the next section we shall take a look at the case with nonzero $\mu_{B}$, where the realization of the supercooling chiral phase transition, as has been discussed in the literature Shao:2024dxt ; Shao:2024ygm , will be recapped in terms of the class of the soft-scale breaking and the Debye mass effect.